Coca-Cola uses a process cost accounting system and a weighted-average cost flow assumption. The department adds...

  1. Coca-Cola uses a process cost accounting system and a weighted-average cost flow assumption. The department adds materials at the beginning of the process and incurs conversion costs uniformly throughout the process. During July, $190,000 of materials costs and $137,100 in conversion costs were charged to the department. The beginning work in process inventory was $93,000 on July 1, comprised of $80,000 of materials costs and $13,000 of conversion costs. (20 pts)

Other data for the month of July are as follows:

      Beginning work in process inventory, 7/1                             25,000 units     (40% complete)

      Units completed and transferred out                                     70,000 units

      Ending work in process inventory, 7/31                                30,000 units     (30% complete)

Instructions

Answer the following questions and show computations to support your answers.

1.   How many physical units have to be accounted for in July?

2.   What are the equivalent units of production for materials and for conversion costs for the month of July?

3.   What is the total cost assigned to the 70,000 units that were transferred out of the process in July?

4.         What is the total cost of the July 31 inventory?
Answer #1

1.

Physical units to be accounted for = Units completed + Ending work in process

= 70,000 + 30,000

= 100,000

2.

Equivalent units = Units completed + Ending work in process X (% of completion)

Material = 70,000 + 30,000 = 100,000

Conversion costs = 70,000 + (30,000 X 30%) = 79,000

3.

Cost per equivalent unit = Total cost / Equivalent units

Material = ($80,000 + $190,000) / 100,000 = $2.70

Conversion costs = ($13,000 + $137,100) / 79,000 = $1.90

Total cost = ($2.70 + $1.90) X 70,000 = $322,000

4.

Cost of July 31st inventory = (30,000 X $2.70) + (30,000 X 30% X $1.90) = $98,100